at CNG gas station, on average 3 CNG operated buses arrive every 15 minutes. what kind of probability distribution would be suitable to describe the number of buses arriving? Also , find the probability that during a sample 15 minutes confirming to this distribution the probability of 2 buses would arrive at the gas station
Answer #1

Mean = 3 buses every 15 minutes

This is a Poisson distribution.

mu = 3

P(X = x) = e-3 * 3x / x!

P(X = 2) = e_3 * 32 / 2! = 0.0747 (ans)